4. After primers have been designed and ordered, they are received in very small quantities as a dry solid and need to be res the value that is (check your discussion forum for your value) is 37.1 nmol.
~0.0371umol.
Answer #1

We have 37.1 nmoles of dried primer. We need to make a 10 \mu M solution.

A 10 \mu M solution has 10 micromoles in 1 litre of solution. However, we only have 37.1 nanomoles, therefore, we need to calculate how much water is to be added to make a 10 micromolar solution with 37.1 nanomoles.

Now,

Molarity = No. of Moles/Volume (in Litre)
10 micromolar = 37.1 nanomoles/x L
=> x Litre = 37.1 nanomoles/10 micromolar
=> x = 37.1 * 10-3 micromoles/10 micromolar = 3.71 * 10-3 litre

=> x = 3.71 mL

To make a 10 micromolar solution with 37.1 nmoles of primer, we need to add 3.71 mL of pure water.